# Understanding differences in applying DETR to natural and medical images

### v1

@@ 3.1 Design Choices
In this section, we outline key design choices of Deformable DETR relevant to the unique characteristics of medical images: input resolution, the number of encoder layers, multi-scale feature fusion, the number of object queries, and two techniques enhancing the decoding process, query initialization and iterative bounding box refinement (IBBR). We investigate if incorporating these elements into model design yields improvements in detection performance on medical imaging data to the same degree they do on natural image datasets.
@@
Downsampling input images is a standard practice in detection models for computational efficiency and to satisfy memory constraints. Natural images can be significantly downsized without losing important features such as edges, shapes, and textures necessary for accurate predictions. Data from many medical imaging modalities, such as X-ray images, CT scans, and whole slide images, are at least an order of magnitude larger. These high-resolution medical images contain fine-grained details such as small lesions or slight changes in tissue density that are crucial for accurate diagnosis (Sabottke and Spieler, 2020; Thambawita et al., 2021). On the other hand, processing high-resolution medical images without any downsampling is often infeasible due to the high computational requirements. Hence, we conduct experiments on medical images with resolution ratios ranging from 25% to 100% to examine the optimal input resolution that balances model accuracy with computational efficiency and memory usage.
@@
Medical image datasets have three distinct characteristics. Firstly, they are typically smaller compared to natural image datasets due to the limited number of patients with specific conditions. Secondly, medical images within one dataset are usually very homogeneous, depicting single body part such as the brain, breast, or chest, with uniform textures and grayscale. Furthermore, while natural images contain hundreds or thousands of objects classes, medical image datasets usually have much fewer classes. For example, NIH Chest X-ray (14 disease classes) (Wang et al., 2017), DDSM (Heath et al., 2001) (2 classes), BraTs (4 classes) (Moawad et al., 2023). This means that, overall, there is much less variation in the data that the network has to model. Given the broadly accepted idea that model complexity should match task complexity (Geman et al., 1992), we suspect that simpler, shallower architectures might be better suited to these medical imaging datasets to mitigate overfitting and enhance training efficiency. Additionally, objects in medical images tend to have more uniform sizes compared to those in natural images. For instance, the standard deviation of standardized object sizes [^1] Standardized object size refers to the area of the bounding box of an object divided by the area of the image, providing a normalized measure of object size relative to the image size. in the MS COCO dataset is 0.16, whereas it is only 0.025 in the NYU Breast Cancer Screening Dataset. This observation leads us to question the necessity of employing multi-scale feature fusion for medical images, which is more beneficial for detecting objects of varying sizes. To explore these hypotheses, we experimented with modifications to the encoder of Deformable DETR, including reducing the number of encoder layers and utilizing fewer scales of feature maps from the backbone.
@@
In DETR, each object query is individually decoded into a bounding box prediction. Thus, the total number of object queries determines the number of bounding boxes predicted per image (Carion et al., 2020). Most DETR models are optimized for natural image datasets such as MS COCO, where a single image can contain up to $100$ objects. Consequently, the number of object queries is usually set to $300$ in DETR models. However, a single medical image typically contains fewer than $10$ objects, with the majority containing only $1$ or no objects at all. For this reason, the hyperparameters suggested in DETR research might not be suitable for medical images, making it crucial to adjust this number for medical image datasets.
@@
Object queries initialization and iterative bounding box refinement (IBBR) are two techniques employed by many DETR models to facilitate the decoding process of the object queries (Zhu et al., 2020; Zhang et al., 2022; Yao et al., 2021). These methods have proven effective in boosting detection performance on natural image datasets, increasing AP by 2.4 on the MS COCO dataset (Zhu et al., 2020). This study aims to evaluate their suitability for medical data. We tested three initialization strategies for the positional and content embeddings of object queries, as characterized by Zhang et al. (2022).
  • Static queries: Both positional and content embeddings are initialized as random learnable embeddings. This generic starting point requires the model to learn from scratch where objects are likely located and what features represent those objects. While offering maximum flexibility, this method may lead to slower learning as the model has to discover spatial and content relationships without prior knowledge. The standard Deformable DETR uses this approach.
  • Pure query selection: Both content and positional embeddings are initialized with selected encoder features. To choose $K$ encoder features that most likely contain objects, the model either uses a regional proposal network (Yao et al., 2021; Chen et al., 2022b) or applies the prediction head to encoder features (Zhu et al., 2020). Our experiments employ the latter approach. This method leverages existing knowledge encoded in the encoder features, guiding the model more directly toward accurate object detection and significantly speeding up the learning process.
  • Mixed query selection: This method combines the previous two methods. By using pure query selection for positional embeddings and static embeddings for content, the model is informed about the likely positions of objects through the spatial priors while retaining flexibility in learning content representations from scratch. DINO (Zhang et al., 2022) suggests that mixed query selection methods provide the best model performance among all three methods.
IBBR, first introduced in Deformable DETR, iteratively updates the reference points of object queries towards the objects of interest in each image. Reference points guide the model’s deformable attention mechanism, indicating which regions to search for objects. These reference points are distributed randomly across the image, ensuring a broad coverage without any prior knowledge about where objects might be located. With IBBR, these reference points can move progressively towards the objects through each decoder layer, providing more accurate signals for the attention. This technique has been applied extensively in DETR family (Zhu et al., 2020; Chen et al., 2022b; Wang et al., 2022; Liu et al., 2022) and has shown to effectively speed up training and improve detection performance. Detailed formalization of this method is in Appendix C.
@@ 3.2 Data and task
We conduct experiments on the NYU Breast Cancer Screening Dataset (Wu et al., 2019), representative for the characteristics of medical image datasets. This dataset contains $229,426$ digital screening mammography exams from $141,472$ patients screened at NYU Langone Health. Each exam includes a minimum of four images, each with a resolution of $2944\times 1920$, covering two standard screening views: craniocaudal (CC) and mediolateral oblique (MLO), for both the left and right breasts. An example of a mammography exam is shown in Figure 3. The dataset is annotated with breast-level cancer labels indicating biopsy-confirmed benign or malignant findings. Moreover, the dataset also provides bounding box annotations, and class labels (benign or malignant) of each visible positive findings. The entire dataset contains $985$ breasts with malignant findings and $5,556$ breasts with benign findings. The dataset is divided into training ($82\%$), validation ($5\%$) and test sets ($13\%$) ensuring a proportional distribution of both benign and malignant cases across the subsets.
Figure 3: An example screening mammography exam. From left to right: left craniocaudal view, left mediolateral oblique view, right craniocaudal view, right mediolateral oblique view. This patient has a benign lesion in the left breast. It is marked with a red bounding box from both views of the left breast.
@@ 3.3 Evaluation Metrics
In this study, we focus on evaluating the models’ ability to detect malignant lesions. We use Average Precision (AP) (Everingham et al., 2010) and the Free-Response Receiver Operating Characteristic curve area (FAUC) (Bandos et al., 2009), which is a frequently used metric in medical image analysis (Yu et al., 2022; Wang et al., 2018; Petrick et al., 2013). Specifically, we focus on FAUC at the rate of 1 false positive per image, referred to as $\text{FAUC}^{1}$, in line with the approach described by Bandos et al. (2009). Following prior works (Jailin et al., 2023; Kolchev et al., 2022; Konz et al., 2023), we define a positive bounding box to have at least 10% 50% Intersection over Union (IoU) with a ground truth box. These thresholds are deemed more appropriate for accurately detecting small-sized objects, such as cancerous lesions. $\mathrm{FAUC}^{1}$ at 0.1 Iou threhold is denoted as $\text{FAUC}^{1}_{10}$. Following the notation of COCO evaluation metrics (Lin et al., 2014), we denote AP at 0.1 IoU threshold as $\text{AP}_{10}$. Additionally, we report the average AP across IoU thresholds ranging from 0.1 to 0.5, in steps size of 0.05, denoted as $\text{AP}_{10,50}$.
@@
  • Localization refers to the task of accurately drawing a bounding box around each ground-truth object. To be consistent with the definition of FAUC and AP, an object is considered successfully localized if the model produces a bounding box overlapping the ground truth box by more than 10% IoU. To quantify localization accuracy, we compute the percentage of ground-truth objects successfully detected by the model. Assume there are $m$ ground-truth objects $G_{i}$ where $i=1,\ldots,m$ and $p$ predicted boxes $P_{j}$ where $j=1,\ldots,p$ in an image. The maximum IoU for a ground-truth bound box $G_{i}$ among all the predicted bounding boxes $P_{j}$ is $\max_{j}(\mathrm{IoU}(P_{j},G_{i}))$. Localization accuracy $L$ is then expressed as
    $L=\frac{1}{m}\sum^{m}_{i=1}\mathds{1}_{\max_{j}\left({\mathrm{IoU}\left(P_{j},G_{i}\right)}\right)\geq 0.1}.$ (1)
@@
    $L_{\mathrm{top10}}=\frac{1}{m}\sum^{m}_{i=1}\mathds{1}_{\max_{j\in S}({\mathrm{IoU}(P_{j},G_{i})})\geq 0.1}.$ (2)
@@ 3.4 Experimental Setup
Our baseline model is a Deformable DETR in its default setting, using a Swin-T Transformer (Liu et al., 2021) as the backbone. The backbone is pretrained with a breast cancer classification task on the same dataset (see Appendix D for details). All models are optimized using the AdamW optimizer (Loshchilov and Hutter, 2017) with a step scheduler for $60$ epochs. The step scheduler reduces the learning rate by a factor of $0.1$ for the last $20$ epochs. We tuned the remaining hyperparameters using random search as detailed in Appendix E. For each experiment, we trained $5$ models with different random seeds and we reported their performance mean and standard deviation.
@@ 4 Results
Our experiments with the five design choices, including input resolutions, encoder layer complexity, multi-scale feature fusion, number of object queries, and two decoding techniques, all demonstrate that modifications of the standard Deformable DETR do not align well with the unique characteristics of medical imaging datasets. This misalignment results in unnecessary computational overhead and sub-optimal performance.
@@
Our experiments reveal a positive correlation between input resolution and detection performance up to a certain threshold (Table 1). Specifically, increasing the resolution ratio from a quarter size ($0.25$) to half size ($0.5$) significantly improves the performance, with a $9.8\%$ increase in $\text{FAUC}^{1}_{10}$, $8.6\%$ in $\text{AP}_{10}$ and $6.4\%$ in $\text{AP}_{10,50}$. Further increasing the resolution to $0.75$ continues to boost performance, though with a smaller margin. However, using full-resolution images results in a performance drop across all metrics. In high-resolution images, objects of interest can have defining features spread across a large area, requiring the model to integrate clues from a broader area. The deformable attention mechanism only focuses on a selective set of keys centered around the reference points, which may miss necessary information in high resolution images. This phenomenon of diminished performance at very high resolutions has been observed by prior studies (Sabottke and Spieler, 2020; Thambawita et al., 2021). Richter et al. (2021) also challenged the notion that higher resolution always equates to better performance and pointed out that different CNN networks have a preferred input size.
It is also noteworthy that the increase in resolution from quarter to full resolution comes at the cost of ×$\times 15$ increase in GFLOPs. To balance the training efficiency and detection performance, subsequent experiments are conducted for half-resolution images.
Table 1: Standard Deformable DETR performance using different input image resolutions. The full image resolution is $2944\times 1920$. The detection performance is measured by AP and FAUC, as defined in Section 3.3. GFLOPs are reported using one billion floating point operations per second as the unit. Using a 75% scale of full resolution yields the best detection performance.
Image Resolution $\text{FAUC}^{1}_{10}$ $\text{AP}_{10}$ $\text{AP}_{10,50}$ GFLOPs
$1.0$ $0.676\pm 0.010$ $0.671\pm 0.005$ $0.477\pm 0.006$ $4367$
$0.75$ $0.688\pm 0.010$ $0.683\pm 0.007$ $0.490\pm 0.007$ $2448$
$0.5$ $0.689\pm 0.008$ $0.669\pm 0.012$ $0.464\pm 0.019$ $1102$
$0.25$ $0.591\pm 0.016$ $0.583\pm 0.007$ $0.400\pm 0.021$ $284$
@@
In this set of experiments, we examined the impact of the number of encoder layers and the necessity of muilti-scale feature fusion in the Deformable DETR architecture.
Table 2: Varying the number of encoder layers in Deformable DETR. The standard Deformable DETR has $6$ encoder layers. For both backbones, we do not observe any significant performance drop when using fewer encoder layers. A small performance decrease is observed for a model with no encoder at all.
backbone #encoder $\text{FAUC}^{1}_{10}$ $\text{AP}_{10}$ $\text{AP}_{10,50}$params GFLOPs
layers
ResNet50 00 $0.643\pm 0.020$ $0.619\pm 0.010$ $0.421\pm 0.025$ $35.4$ $536$
ResNet50 $1$ $0.656\pm 0.011$ $0.624\pm 0.016$ $0.439\pm 0.005$ $36.2$ $624$
ResNet50 $3$ $0.655\pm 0.016$ $0.627\pm 0.013$ $0.439\pm 0.015$ $37.7$ $801$
ResNet50 $6$ $0.657\pm 0.009$ $0.626\pm 0.009$ $0.436\pm 0.011$ $40.0$ $1067$
Swin-T 00 $0.681\pm 0.013$ $0.662\pm 0.013$ $0.458\pm 0.023$ $35.9$ $570$
Swin-T $1$ $0.684\pm 0.009$ $0.672\pm 0.011$ $0.463\pm 0.014$ $36.7$ $659$
Swin-T $3$ $0.688\pm 0.012$ $0.677\pm 0.011$ $0.470\pm 0.019$ $38.3$ $836$
Swin-T $6$ $0.689\pm 0.008$ $0.669\pm 0.012$ $0.464\pm 0.019$ $40.5$ $1102$
As the backbone critically affects feature extraction, to reach more generalizable conclusions, we conducted this experiment with two distinct backbones, ResNet50 and Swin-T. For both backbones, models with merely $1$ and $3$ encoder layers matched the performance of models with $6$ encoder layers across all three detection metrics (Table 2). Impressively, this reduction in the number of layers resulted in up to $40\%$ decrease in GFLOPs. Furthermore, when using Swin-T as the backbone, the encoder-free models achieve nearly the same detection performance as their 6-layer counterparts, with less than a $1\%$ drop across all three detection metrics, while cutting the GFLOPs almost by half.

### v2

@@ 3.1 Design Choices
In this section, we outline key design choices in Deformable DETR that are relevant to the unique characteristics of medical images: input resolution, the number of encoder layers, multi-scale feature fusion, the number of object queries, and two techniques enhancing the decoding process, query initialization and iterative bounding box refinement (IBBR). We investigated whether these components, which improve performance on natural image datasets, offer similar benefits when applied to medical imaging tasks.
@@
Downsampling is commonly used in detection models to reduce computational cost and satisfy memory constraints. Natural images can be significantly downsized to $224\times 224$ or $256\times 256$ pixels without losing important features, such as edges, shapes, and textures that are necessary for accurate predictions. In contrast, medical images are often an order of magnitude larger. For example, X-ray images can reach up to $2500\times 3056$ pixels (Johnson et al., 2019), and CT scans are typically $512\times 512$ pixels (Setio et al., 2017). These high-resolution medical images contain fine-grained details, such as small lesions or slight changes in tissue density, which are crucial for an accurate diagnosis (Sabottke and Spieler, 2020; Thambawita et al., 2021). However, processing high-resolution medical images is often infeasible due to the high computational requirements. To address this trade-off, we evaluate performance across input resolutions ranging from 25% to 100% of the original size, aiming to identify the optimal input resolution that balances model accuracy with computational efficiency and memory usage.
@@
Medical imaging datasets differ from natural image datasets in several important ways. First, they are typically smaller due to limited patient availability. Second, images within a dataset tend to be homogeneous, focusing on a single body part, such as the brain, breast, or chest, with uniform grayscale textures (Figure 2). Third, while natural images contain hundreds or thousands of object classes, medical image datasets usually have far fewer object classes. For example, NIH Chest X-ray contains 14 classes (Wang et al., 2017), DDSM has 2 (Heath et al., 2001), and BraTS has 4 (Moawad et al., 2023). As a result, there is much less variation in the data that the model has to capture. Given the principle that model complexity should align with task complexity (Geman et al., 1992), we suspect that simpler, shallower architectures might be more appropriate for medical image analysis, helping mitigate overfitting and improve training efficiency. In addition, object sizes in medical images are typically more uniform than in natural scenes. For example, the standard deviation of normalized object sizes [^1] Normalized object size refers to the area of the bounding box divided by the total image area. is 0.025 in the NYU Breast Cancer Screening Dataset and 0.001 in LUNA16, compared to 0.16 in MS COCO. This raises questions about the use of multi-scale feature fusion in this domain, a technique primarily intended to improve detection across diverse object sizes. To investigate these hypotheses, we experimented with modifications to the encoder of Deformable DETR, including reducing the number of encoder layers and utilizing fewer scales of feature maps from the backbone.
@@
In DETR, each object query is individually decoded into a bounding box prediction. Thus, the total number of object queries determines how many objects the model can detect per image (Carion et al., 2020). Most DETR models are optimized for natural image datasets such as MS COCO, where a single image can contain up to $100$ objects. Consequently, the number of object queries is usually set to $300$ in DETR models. In contrast, medical images rarely contain more than 10 objects, and most have only one or none. As a result, the default number of object queries used in standard DETR implementations may be excessive for medical applications, potentially leading to unnecessary computation or degraded performance. We therefore examine how reducing the number of object queries affects detection accuracy and efficiency on medical image datasets.
@@
Many DETR variants apply object queries initialization and iterative bounding box refinement (IBBR) to improve query decoding and increase detection accuracy (Zhu et al., 2020; Zhang et al., 2022; Yao et al., 2021). These methods have proven effective in boosting detection performance on natural image datasets, increasing average precision by 2.4 on the MS COCO dataset (Zhu et al., 2020). In this study, we evaluate their effectiveness in the medical imaging domain. We tested three initialization strategies for the positional and content embeddings of object queries, as characterized by Zhang et al. (2022).
  • Static queries Both positional and content embeddings are randomly initialized as learnable embeddings. This offers maximum flexibility, but requires the model to learn where objects are likely located and what features represent those objects from scratch, potentially slowing convergence. Standard Deformable DETR uses this approach.
  • Pure query selection Both content and positional embeddings are initialized from selected encoder features. Following Zhu et al. (2020), we apply the prediction head to the encoder output to select the top-$K$ features. Some other works use a regional proposal network (Yao et al., 2021; Chen et al., 2022b). This leverages encoder knowledge to guide object queries and significantly accelerates training.
  • Mixed Query Selection: Positional embeddings are initialized from encoder features (as above), while content embeddings remain randomly initialized. This hybrid strategy informs about the likely positions of objects through spatial priors while retaining flexibility in learning content representations from scratch. DETR with Improved DeNoising Anchor Boxes (DINO) (Zhang et al., 2022) found that this method yields the best performance.
IBBR, first introduced in Deformable DETR, iteratively updates the reference points of object queries towards the objects of interest in each image. These reference points guide the deformable attention toward relevant regions to search for objects. Initially, they are randomly distributed across the image, ensuring broad coverage without any prior knowledge about where objects might be located. With IBBR, these reference points can move progressively towards the objects through each decoder layer, providing more accurate signals for attention. This technique has been extensively applied in subsequent DETR variants (Zhu et al., 2020; Chen et al., 2022b; Wang et al., 2022; Liu et al., 2022) and has been shown to effectively speed up training and improve detection performance. A detailed formalization is provided in the Appendix C.
@@ 3.2 Data and task
NYU Breast Cancer Screening Dataset (NYU Breast)
(Wu et al., 2019) contains $229,426$ digital screening mammography exams from $141,472$ patients screened at NYU Langone Health. Each exam includes a minimum of four images, each with a resolution of $2944\times 1920$, covering two standard screening views: craniocaudal (CC) and mediolateral oblique (MLO), for both the left and right breasts. An example of a mammography exam is shown in Figure 4. The dataset is annotated with breast-level cancer labels indicating biopsy-confirmed benign or malignant findings. Moreover, the dataset also provides bounding box annotations, and class labels (benign or malignant) of each visible positive findings. The entire dataset contains $985$ breasts with malignant findings and $5,556$ breasts with benign findings. The dataset is divided into training ($82\%$), validation ($5\%$) and test sets ($13\%$) ensuring a proportional distribution of benign and malignant cases across the subsets.
LUNA16
LUNA16 (Setio et al., 2017) is a public chest CT dataset for lung nodule detection, containing 888 3D chest CT scans with annotated nodule locations. We selected this dataset because it exemplifies key characteristics of medical imaging (Figure 5): (1) high resolution (typically 512 x 512 pixels per slice), necessary for capturing fine details; (2) small objects of interest, as nodules are subtle and occupy only a small portion of each scan; and (3) class imbalance, as nodules are relatively rare. Each nodule is annotated in 3D with a center point $(x,y,z)$ and diameter. To convert them into 2D bounding boxes, we identify the slices intersecting each nodule along the z-coordinate and project the center point to 2D $(x,y)$ coordinates on each slice. Using the diameter, we define a 2D bounding box around this point, allowing slice-by-slice nodule detection. Since DETR is designed for 2D object detection, we treat each 2D slice as an independent input to the model, enabling nodule detection in each slice separately. The dataset is randomly split for training (666 scans, 75%), validation (88 scans, 10%), and test (134 scans, 15%).
Figure 4: An example screening mammography exam. From left to right: left craniocaudal view (L-CC), left mediolateral oblique view (L-MLO), right craniocaudal view (R-CC), right mediolateral oblique view (R-MLO). This patient has a benign lesion in the left breast. It is marked with a red bounding box from both views of the left breast.
Figure 5: 2D CT scans with nodule annotations of LUNA16 Each image shows a single axial CT slice with a red bounding box indicating the location of a lung nodule. These examples illustrate the small size and subtle appearance of nodules, highlighting the challenges of object detection in medical imaging.
@@ 3.3 Evaluation Metrics
In this study, we focus on evaluating the ability of the models to detect malignant lesions. We use Average Precision (AP) (Everingham et al., 2010) and the Free-Response Receiver Operating Characteristic curve area (FAUC) (Bandos et al., 2009), which is a frequently used metric in medical image analysis (Yu et al., 2022; Wang et al., 2018; Petrick et al., 2013). Specifically, we focus on FAUC at the rate of 1 false positive per image or smaller, referred to as $\mathrm{FAUC}^{1}$, in line with the approach described by Bandos et al. (2009). To formalize $\mathrm{FAUC}^{1}$, we introduce the following notation:
Let N denote the total number of images, indexed by $n=1,2,..,N$. Each image $n$ contains $L_{n}$ lesions with $L_{total}=\sum^{N}_{1}L_{n}$ being the total lesions in the dataset. For each image $n$, a detection model produces a set of candidate detections $\mathcal{D}_{n}=\{d_{n}^{1},d_{n}^{2},...\}$, each with a confidence score $s(d_{n}^{j})$. By varying a decision threshold $\tau$, one can include only those detections whose scores exceed $\tau$, denoted $\mathcal{D}_{n}(\tau)=\{d\in\mathcal{D}_{n}:s(d)\leq\tau\}$.
Following prior works (Jailin et al., 2023; Kolchev et al., 2022; Konz et al., 2023), we define a positive bounding box to have at least 10% Intersection over Union (IoU) with a ground truth box. These thresholds are deemed more appropriate for accurately detecting small-sized objects, such as cancerous lesions. The $\mathrm{FAUC}^{1}_{10}$ metric integrates the true positive rate (TPR) over false positives per image (FPI) from 0 to 1:
$\mathrm{FAUC}^{1}_{10}=\int^{1}_{0}\frac{1}{L_{total}}\sum^{n}_{n=1}\sum^{L_{n}}_{m=1}1(\exists d\in\mathcal{D}_{n}(\tau(u)):d\leftrightarrow L_{n}^{m})du,$
where:
  • $\tau(u)$ is the threshold achieving $\mathrm{FPI}=\frac{1}{N}\sum_{N=1}^{N}|\{d\in\mathcal{D}_{n}(\tau):d\nleftrightarrow L_{n}^{m},\forall m\in L_{n}\}=u$
  • $d\leftrightarrow L_{n}^{m}$ indicates the lesion $m$ is detected by the prediction $d$ in image $n$
Following the notation of integrated average precision in PASCAL VOC 2012  (Salton, 1983; Everingham et al., 2010), we denote AP at 0.1 IoU threshold as $\mathrm{AP}_{10}$. Additionally, we report the average AP across IoU thresholds ranging from 0.1 to 0.5, in steps size of 0.05, denoted as $\mathrm{AP}_{10,50}$.
@@
  • Localization refers to the task of accurately drawing a bounding box around each ground-truth object. To be consistent with the definition of FAUC and AP, an object is considered successfully localized if the model produces a bounding box overlapping the ground truth box by more than 10% IoU. To quantify localization accuracy, we compute the percentage of ground-truth objects successfully detected by the model. Assume there are $m$ ground-truth objects $G_{i}$ where $i=1,\ldots,m$ and $p$ predicted boxes $P_{j}$ where $j=1,\ldots,p$ in an image. The maximum IoU for a ground truth bound box $G_{i}$ among all predicted bounding boxes $P_{j}$ is $\max_{j}(\mathrm{IoU}(P_{j},G_{i}))$. Localization accuracy $L$ is then expressed as
    $L=\frac{1}{m}\sum^{m}_{i=1}\mathds{1}_{\max_{j}\left({\mathrm{IoU}\left(P_{j},G_{i}\right)}\right)\geq 0.1},$ (1)
    where the indicator function is defined as
    $\mathds{1}_{\max_{j}\left({\mathrm{IoU}\left(P_{j},G_{i}\right)}\right)\geq 0.1}=\begin{cases}1,&\mathrm{if}\max_{j}\left(\mathrm{IoU}\left(P_{j},G_{i}\right)\right)\geq 0.1\\ 0,&\mathrm{otherwise}.\end{cases}$
@@
    $L_{\mathrm{top10}}=\frac{1}{m}\sum^{m}_{i=1}\mathds{1}_{\max_{j\in S}({\mathrm{IoU}(P_{j},G_{i})})\geq 0.1},$ (2)
    where the indicator function is defined as
    $\mathds{1}_{\max_{j\in S}\left({\mathrm{IoU}\left(P_{j},G_{i}\right)}\right)\geq 0.1}$ $=$
    $\begin{cases}1,&\mathrm{if}\max_{j\in S}\left(\mathrm{IoU}\left(P_{j},G_{i}\right)\right)\geq 0.1\\ 0,&\mathrm{otherwise}.\end{cases}$
@@ 3.4 Experimental Setup
Our baseline model is Deformable DETR in its default setting, using a Swin-T backbone (Liu et al., 2021). For the NYU Breast Cancer Screening dataset, the backbone is pretrained on a breast cancer classification task using the same dataset (see Appendix D for details). Models are trained for 60 epochs on NYU Breast and 100 epochs on LUNA16. We used a batch size of 2 for NYU Breast and 32 for LUNA16. All models use the AdamW optimizer (Loshchilov and Hutter, 2017) with a step learning rate scheduler, which reduces the learning rate by a factor of 0.1 during the final 20 epochs. We tuned the hyperparameters using random search as detailed in Appendix E. To account for training variability, we train five models with different random seeds for each experiment and report the mean and standard deviation of their performance. All training is conducted using a single NVIDIA A100 GPU.
@@ 4 Results
Our experiments across the five design choices, including input resolutions, encoder layer complexity, multi-scale feature fusion, number of object queries, and two decoding techniques, reveal that standard Deformable DETR configurations do not align well with the unique characteristics of medical imaging datasets. This misalignment results in unnecessary computational overhead and sub-optimal performance.
@@
Our experiments reveal a positive correlation between input resolution and detection performance, up to a certain point, for both the NYU Breast and LUNA16 datasets (Table 1). Specifically, increasing the resolution from 25% to 50% of the original image size significantly improves performance. On NYU Breast, this yields gains of 9.8% in $\mathrm{FAUC}^{1}_{10}$, 8.6% in $\mathrm{AP}_{10}$, and 6.4% in $\mathrm{AP}_{10,50}$. Similarly, LUNA16 shows improvements of 5.9%, 11.8%, and 6.0% in the corresponding metrics. Raising the resolution to $75\%$ continues to improve performance, although with diminishing returns in the NYU Breast. Interestingly, full-resolution images result in a decline in performance across all metrics on both datasets. This may be attributed to the limitations of the deformable attention mechanism. In high-resolution images, objects of interest may be distributed across a wider spatial area. The deformable attention mechanism only focuses on a selective set of keys centered around the reference points, which may miss necessary information in high resolution images. This phenomenon aligns with previous findings that question the assumption that higher resolution always improves performance (Sabottke and Spieler, 2020; Thambawita et al., 2021; Richter et al., 2021).
It is also important to note the computational trade-off: increasing the input resolution from 25% to 100% results in a 10–15$\times$ increase in GFLOPs. To balance accuracy with computational efficiency, we used half-resolution images (50%) for subsequent NYU Breast experiments and 75% resolution for LUNA16, as these settings offer the best trade-off between performance and resource usage.
Table 1: Standard Deformable DETR performance using different input image resolutions. The full resolution of images from NYU Breast is $2944\times 1920$ and LUNA16 is $512\times 512$. The detection performance is measured by AP and FAUC, as defined in Section 3.3. GFLOPs are reported using one billion floating point operations per second as the unit.
Dataset Image $\mathrm{FAUC}^{1}_{10}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pm\mathrm{SD}}$ $\mathrm{AP}_{10}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pm\mathrm{SD}}$ $\mathrm{AP}_{10,50}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pm\mathrm{SD}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L\pm\mathrm{SD}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L_{top10}\pm\mathrm{SD}}$ GFLOPs Resolution NYU $1.0$ $0.676\pm 0.010$ $0.671\pm 0.005$ $0.477\pm 0.006$ $0.925\pm 0.004$ $0.843\pm 0.003$ $4367$ $0.75$ $0.688\pm 0.010$ $0.683\pm 0.007$ $0.490\pm 0.007$ $0.930\pm 0.004$ $0.862\pm 0.005$ $2448$ $0.5$ $0.689\pm 0.008$ $0.669\pm 0.012$ $0.464\pm 0.019$ $0.920\pm 0.005$ $0.856\pm 0.009$ $1102$ $0.25$ $0.591\pm 0.016$ $0.583\pm 0.007$ $0.400\pm 0.021$ $0.897\pm 0.003$ $0.799\pm 0.01$ $284$ LUNA $1.0$ $0.521\pm 0.011$ $0.363\pm 0.018$ $0.221\pm 0.015$ $0.959\pm 0.006$ $0.651\pm 0.012$ $3425$ $0.75$ $0.537\pm 0.007$ $0.392\pm 0.013$ $0.295\pm 0.005$ $0.959\pm 0.002$ $0.663\pm 0.011$ $1966$ $0.5$ $0.488\pm 0.003$ $0.340\pm 0.014$ $0.221\pm 0.019$ $0.941\pm 0.011$ $0.628\pm 0.011$ $986$ $0.25$ $0.461\pm 0.030$ $0.304\pm 0.011$ $0.161\pm 0.022$ $0.951\pm 0.014$ $0.604\pm 0.018$ $340$
@@
We investigated the effect of varying the number of encoder layers in Deformable DETR and evaluated whether the full encoder depth is necessary for medical imaging tasks. To ensure generalizability, we conducted experiments using two distinct backbones, ResNet50 and Swin-T.
On the NYU Breast dataset, for both backbones, reducing encoder layers from six to one or three results in comparable performance in all three detection metrics, while reducing GFLOPs by up to 40% (Table 2). In particular, encoder-free models (0 layers) with Swin-T maintain performance within 1% of the full 6-layer model, while cutting computation nearly half. This is likely because Swin-T was pretrained on the same mammography dataset, allowing it to extract strong task-specific features and reducing its reliance on the encoder. On LUNA16, we observed a similar pattern, with one or three encoder layers yielding a performance comparable to that of the full model, but the encoder-free models fail completely. This is likely due to not pre-training Swin-T backbone on lung CT, highlighting that when the backbone is not adapted to the target domain, some encoder capacity becomes necessary. Nevertheless, even with minimal encoder depth (one layer), the model achieved strong results while significantly lowering computational cost, from 1966 to 1225 GFLOPs.